OpenStudy (anonymous):

a mass m moving with velocity v0 along a line y=-a and away from the origin find the magnitude of angular momentum with respect to origin

OpenStudy (anonymous):

please answer me fast

OpenStudy (anonymous):

Angular momentum is calculated as \[L = R \times P \] where R is the position vector and P is the momentum!.. when you want magnitude \[\left| L \right| = rp \sin \theta \] where theta is angle between the position vector and the linear momentum!.. so can you calculate now?

OpenStudy (anonymous):